OpenStudy (anonymous):

Unit 6 Rational Expression Simplify u^2 - 2u u^2 - 4 A. u u + 4 U doesnt equal -2 B. 1 u + 2 U doesnt equal c. u u + 2 u doesnt -2 and u doesnt 2 D. u-2 u + 2 U doesnt equal -2 and U doesnt equal 2

OpenStudy (anonymous):

\[\frac{ u^2-2u }{ u^2-4 }\]

OpenStudy (anonymous):

For this one i'll have to factor these hold on

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{ u^2-2u }{ (u-2)(u+2) }\]

OpenStudy (triciaal):

@Schen421 also factor the top

OpenStudy (anonymous):

\[\frac{ (u)(u-2) }{ (u+2)(u-2) }\]

OpenStudy (anonymous):

Alright so since one of the numerator and denominator are the same, you can cancel out the (u-2) from the top and bottom

OpenStudy (anonymous):

Remember that the x values for the denominator cannot be any number that makes the denominator a 0! if it makes the denominator a 0 then x cannot be that number

OpenStudy (anonymous):

Okay so the (\[\frac{ u}{ u + 2} u \neq -2 and u \neq 2\]

OpenStudy (anonymous):

Yes! Nicely done! =)

OpenStudy (anonymous):

thanks